A Figurines Company sells figurines to fan. The university buys figurines from the manufacturer at $6/unit. They send orders to the manufacturer, costing $50/order and there is an average lead time of twenty days for the orders to arrive from the manufacturer. Their inventory carrying cost is 18% of the unit cost. The average daily demand is 30 figurines per day. They are open for business 275 days per year.please show work

            a. How many figurines should the university order each time?

            b. How many orders will they make per year?

            c. Compute average inventory

            d. What is the annual ordering cost?

            e. What is the annual inventory carrying cost?
Answer #1

Answer: S = Setup costs (order cost) = $50
D = Demand (quantity demand per year) = average daily demand*working days in a year= 30*275= 11,250

Figurines cost= $6 per unit

H = Holding costs = 18% of unit cost= 0.18*6= $1.08 per year

  1. Firm should order according to EOQ, EOQ = √(2SD / H) = 1020.6
  2. Total no. of orders= total demand/EOQ= 11.02
  3. Average inventory= EOQ/2= 510.3
  4. total annual order cost= no. of orders *order cost= 11.02*50= $551
  5. total annual holding cost= (EOQ/2)*holding cost= (510.3)*1.08= $551

Please ask, if you have any doubts through the comment section. Do rate the answer